# Many-hypercube codes: High-rate quantum error-correcting codes for high-performance fault-tolerant quantum computation

### v1

@@ Abstract
Conventional approaches to quantum error correction for fault-tolerant quantum computation are based on encoding a single logical qubit into many physical qubits, resulting in asymptotically zero encoding rates and therefore huge resource overheads. To overcome this issue, high-rate quantum codes, such as quantum low-density parity-check codes, have been studied over the past decade. However, such codes have complex structure, making it difficult to perform logical gate operations in parallel without sacrificing their advantage. Observing the simple structure and high rates of quantum error-detecting codes, here we propose concatenated high-rate quantum error-detecting codes as a new family of high-rate quantum codes. Their simple structure allows for a geometrical interpretation using hypercubes, each of which corresponds to a logical qubit. We thus call them many-hypercube codes. The encoding rate is remarkably high, e.g., 30% (64 logical qubits are encoded into 216 physical qubits). Developing a dedicated high-performance decoder, we achieve high error thresholds even in a circuit-level noise model. Logical gate operations are also parallelizable. Thus, the many-hypercube codes will pave the way to high-performance fault-tolerant quantum computation.
@@ I Introduction
In contrast, quantum low-density parity-check (qLDPC) codes [17] are known for their ability to achieve constant encoding rates, and therefore have been studied over the past decade to mitigate the above overhead issue [18, 19]. Various kinds of qLDPC codes have been proposed [20, 21, 22, 23, 24, 25, 26, 27, 28], high-performance decoders for them have also been developed [29, 30, 31, 32, 33, 34, 35], and their physical implementations have recently been proposed [36, 37, 38, 39]. However, they have relatively complex structure, making it difficult to implement logical gate operations in a fault-tolerant manner. A few methods for this purpose have been proposed [18, 37, 38, 39, 40, 41, 42], but parallel execution of logical gates maintaining the advantage of the qLDPC codes, i.e., low overheads is still challenging.
@@
In this work, we propose another family of high-rate concatenated quantum codes. The characteristic feature of our proposal is the use of quantum error-detecting codes [2], which have distance 2 and therefore can detect an error but cannot correct it. By concatenating the error-detecting codes, we can increase the code distance and thus obtain error-correcting codes. The advantage of the quantum error-detecting codes is their simple structure. Harnessing this advantage, Knill proposed concatenated quantum error-detecting codes called the ${C_{4}/C_{6}}$ scheme and achieved very high performance, i.e., the first realization of the error threshold exceeding 1% in a circuit-level noise model [44]. (Recently, the concatenation of the ${C_{4}/C_{6}}$ scheme and the concatenated quantum Hamming codes mentioned above has been proposed to improve the performance of the latter [45], but this still has the above-mentioned two issues, that is, the rate becomes rather lower and the decoding is still based on hard-decision decoding.) However, the ${C_{4}/C_{6}}$ scheme is based on a single-logical-qubit encoding, like the surface code, and therefore its encoding rate approaches zero rapidly as the code size increases. Unlike the ${C_{4}/C_{6}}$ scheme, we concatenate high-rate quantum error-detecting codes. In this work, we focus on the ${[\![6,4,2]\!]}$ code (see Appendix B for its details), where four qubits are encoded into six qubits and the distance is 2. (The reason for choosing this code is its relatively high rate and small size. The use of other codes, such as ${[\![4,2,2]\!]}$ and ${[\![8,6,2]\!]}$, is also interesting, but left for future work.) By concatenating it ${(L-1)}$ times, we obtain the ${[\![6^{L},4^{L},2^{L}]\!]}$ code, which we refer to as level-$L$ C642. Although it is not a constant-rate code, that is, its rate, ${(4/6)^{L}}$, approaches zero as $L$ becomes larger, the rate is remarkably high for small $L$, e.g., 30% (20%) at the level 3 (4) with distance 8 (16), which is higher than not only the surface code, but also well-studied qLDPC codes [36, 37, 38, 39]. Thus the C642, which we also call “many-hypercube codes” for the reason explained later (Fig. 1), will be promising as a near-term target. Note that the number of logical qubits is not limited to $K^{(L)}=4^{L}$. Using M code blocks encoded with the level-$L$ C642 as fault-tolerant quantum registers [43], we can use $MK^{(L)}$ logical qubits for fault-tolerant quantum computing.
@@
We developed a high-performance decoder dedicated to the C642 based on level-by-level minimum distance decoding. (Our proposed decoding method will also be useful for other concatenated codes, such as the above concatenated quantum Hamming codes.) Using this decoder, we achieved a threshold of 5.6% for bit-flip errors, which is comparable to the surface code (10.9%) [46] and a 4%-rate qLDPC (hypergraph product) code (7.5%) [33]. We also propose fault-tolerant zero-state encoders for the C642. Using them, we achieved a threshold of 0.9% for a logical controlled-NOT (CNOT) gate in a circuit-level noise model. Finally, we explain how to perform logical gates for the C642 in parallel.
II Definition of C642 (many-hypercube codes)
@@
Figure 1: Visualization of the structure of the level-3 C642 (many-hypercube code). The vertices correspond to physical qubits. Examples of logical $Z$ and $X$ operators and stabilizers are highlighted by blue ($Z$) and red ($X$). See the main text for the labelling. In general, logical $Z$ and $X$ operators for the level-$L$ C642 correspond to $L$-dimensional hypercubes, the number of which is equal to that of the logical qubits.
@@
Using the ${[\![6,4,2]\!]}$ code recursively, e.g., the level-3 C642 is defined as follows. Each of physical and encoded qubits is labelled by three integers (corresponding to the level) and denoted by Q with subscripts of the integers. The corresponding operators are also labelled similarly. The four level-1 qubits $\{\mathrm{Q}_{i^{\prime},j^{\prime},k}^{(1)}|k=1,\ldots,4\}$ are encoded into the corresponding six physical qubits $\{\mathrm{Q}_{i^{\prime},j^{\prime},k^{\prime}}|k^{\prime}=1,\ldots,6\}$ ($i^{\prime},j^{\prime}=1,\ldots,6$), where the corresponding two stabilizers and logical $Z$ and $X$ operators are defined as follows (see Appendices A and B for the definitions of elementary gates and the ${[\![6,4,2]\!]}$ code, respectively):
$SZ_{i^{\prime},j^{\prime}}^{(1)}$ $=Z_{i^{\prime},j^{\prime},1}Z_{i^{\prime},j^{\prime},2}Z_{i^{\prime},j^{\prime},3}Z_{i^{\prime},j^{\prime},4}Z_{i^{\prime},j^{\prime},5}Z_{i^{\prime},j^{\prime},6},$ (1)
$SX_{i^{\prime},j^{\prime}}^{(1)}$ $=X_{i^{\prime},j^{\prime},1}X_{i^{\prime},j^{\prime},2}X_{i^{\prime},j^{\prime},3}X_{i^{\prime},j^{\prime},4}X_{i^{\prime},j^{\prime},5}X_{i^{\prime},j^{\prime},6},$ (2)
$Z_{i^{\prime},j^{\prime},1}^{(1)}$ $=Z_{i^{\prime},j^{\prime},1}Z_{i^{\prime},j^{\prime},2},~{}X_{i^{\prime},j^{\prime},1}^{(1)}=X_{i^{\prime},j^{\prime},2}X_{i^{\prime},j^{\prime},3},$ (3)
$Z_{i^{\prime},j^{\prime},2}^{(1)}$ $=Z_{i^{\prime},j^{\prime},2}Z_{i^{\prime},j^{\prime},3},~{}X_{i^{\prime},j^{\prime},2}^{(1)}=X_{i^{\prime},j^{\prime},1}X_{i^{\prime},j^{\prime},2},$ (4)
$Z_{i^{\prime},j^{\prime},3}^{(1)}$ $=Z_{i^{\prime},j^{\prime},4}Z_{i^{\prime},j^{\prime},5},~{}X_{i^{\prime},j^{\prime},3}^{(1)}=X_{i^{\prime},j^{\prime},5}X_{i^{\prime},j^{\prime},6},$ (5)
$Z_{i^{\prime},j^{\prime},4}^{(1)}$ $=Z_{i^{\prime},j^{\prime},5}Z_{i^{\prime},j^{\prime},6},~{}X_{i^{\prime},j^{\prime},4}^{(1)}=X_{i^{\prime},j^{\prime},4}X_{i^{\prime},j^{\prime},5}.$ (6)
@@
Similarly, the four level-2 qubits $\{\mathrm{Q}_{i^{\prime},j,k}^{(2)}|j=1,\ldots,4\}$ are encoded into the corresponding six level-1 qubits $\{\mathrm{Q}_{i^{\prime},j^{\prime},k}^{(1)}|j^{\prime}=1,\ldots,6\}$ ($i^{\prime}=1,\ldots,6,k=1,\ldots,4$) and the four level-3 qubits $\{\mathrm{Q}_{i,j,k}^{(3)}|i=1,\ldots,4\}$ are encoded into the corresponding six level-2 qubits $\{\mathrm{Q}_{i^{\prime},j,k}^{(2)}|i^{\prime}=1,\ldots,6\}$ ($j,k=1,\ldots,4$). In this case, we use the 64 level-3 encoded qubits as logical qubits as a fault-tolerant quantum register for quantum computing. In general, the level-$L$ C642 is defined similarly.
@@
Interestingly, the above code structure can be visualized as shown in Fig. 1, where the logical $Z$ and $X$ operators correspond to cubes. In general, the logical $Z$ and $X$ operators of the level-$L$ C642 correspond to $L$-dimensional hypercubes in similar visualization, where the number of hypercubes is equal to that of logical qubits. We thus call the C642 “many-hypercube codes.” [The “many” is used to distinguish it from a conventional hypercube code [47], which is the ${[\![2^{D},D,2]\!]}$ code defined on a single $D$-dimensional hypercube.] The number of the vertices of each $L$-dimensional hypercube is $2^{L}$, which is equal to the code distance. That is, the size of the many-hypercube codes is increased by increasing not the edge length but the dimension while keeping the edge length. In contrast, the logical operators and stabilizers of topological codes, such as the surface code [4, 5, 6, 7, 8, 9, 10] and the color code [48, 49, 50, 51], correspond to the edges and faces, respectively, and their code sizes are increased by increasing the edge length. Unlike the topological codes, the many-hypercube codes require interactions beyond nearest-neighbor ones. This is experimentally challenging, but recent experimental advances in ion-trap [52, 53, 54] and neutral-atom [55, 56, 57, 58] systems are encouraging. In fact, hypercube connectivity has already been realized experimentally [58].
III Decoding of C642

### v3

# Many-hypercube codes: High-rate quantum error-correcting codes for high-performance fault-tolerant quantum computing

@@ Abstract
Standard approaches to quantum error correction for fault-tolerant quantum computing are based on encoding a single logical qubit into many physical ones, resulting in asymptotically zero encoding rates and therefore huge resource overheads. To overcome this issue, high-rate quantum codes, such as quantum low-density parity-check codes, have been studied over the past decade. In this case, however, it is difficult to perform logical gates in parallel while maintaining low overheads. Here we propose concatenated high-rate small-size quantum error-detecting codes as a new family of high-rate quantum codes. Their simple structure allows for a geometrical interpretation using hypercubes corresponding to logical qubits. We thus call them many-hypercube codes. They can realize both high rates, e.g., 30% (64 logical qubits are encoded into 216 physical ones), and parallelizability of logical gates. Developing a high-performance decoder and encoders dedicated to the proposed codes, we achieve high error thresholds even in a circuit-level noise model. Thus, the many-hypercube codes will pave the way to high-performance fault-tolerant quantum computing.
@@ I Introduction
In contrast, quantum low-density parity-check (qLDPC) codes [17] are known for their ability to achieve constant encoding rates, and therefore have been studied over the past decade to mitigate the above overhead issue [18, 19]. Various kinds of qLDPC codes have been proposed [20, 21, 22, 23, 24, 25, 26, 27, 28], high-performance decoders for them have also been developed [29, 30, 31, 32, 33, 34, 35], and their physical implementations have recently been proposed [36, 37, 38, 39, 40]. However, they have relatively complex structure, making it difficult to implement logical gate operations in a fault-tolerant manner. A few methods for this purpose have been proposed [18, 39, 40, 41, 42, 43], but parallel execution of logical gates maintaining the advantage of the qLDPC codes, i.e., low overheads is still challenging.
@@
In this work, we propose another family of high-rate concatenated quantum codes. The characteristic feature of our proposal is the use of quantum error-detecting codes [2], which have distance 2 and therefore can detect an error but cannot correct it. By concatenating the error-detecting codes, we can increase the code distance and thus obtain error-correcting codes. The advantage of the quantum error-detecting codes is their simple structure. Harnessing this advantage, Knill proposed concatenated quantum error-detecting codes called the ${C_{4}/C_{6}}$ scheme and achieved very high performance, i.e., the first realization of the error threshold exceeding 1% in a circuit-level noise model [45]. (Recently, the concatenation of the ${C_{4}/C_{6}}$ scheme and the concatenated quantum Hamming codes mentioned above has been proposed to improve the performance of the latter [46], but this still has the above-mentioned two issues, that is, the rate becomes rather lower and the decoding is still based on hard-decision decoding.) However, the ${C_{4}/C_{6}}$ scheme is based on a single-logical-qubit encoding, like the surface code, and therefore its encoding rate approaches zero rapidly as the code size increases. Unlike the ${C_{4}/C_{6}}$ scheme, the proposed concatenated codes have high encoding rates. In this work, we focus on the ${[\![6,4,2]\!]}$ code, which encodes four qubits into six qubits and has distance 2. (The reason for choosing this code is its relatively high rate and small size. The use of other codes, such as ${[\![4,2,2]\!]}$ and ${[\![8,6,2]\!]}$, is also interesting, but left for future work.) By concatenating it ${(L-1)}$ times, we obtain the ${[\![N^{(L)},K^{(L)},D^{(L)}]\!]}={[\![6^{L},4^{L},2^{L}]\!]}$ code, which we refer to as the ${[\![6^{L},4^{L},2^{L}]\!]}$ level-$L$ many-hypercube code or simply the level-$L$ many-hypercube code for the reason explained later. Although it is not a constant-rate code, that is, its rate, $K^{(L)}/N^{(L)}={(4/6)^{L}}$, approaches zero as $L$ becomes larger, the rate is remarkably high for small $L$, e.g., 30% (20%) at the level 3 (4) with distance 8 (16), which is higher than not only the surface code, but also well-studied qLDPC codes [36, 37, 38, 39, 40]. Thus the many-hypercube codes will be promising as a near-term target. Note that the number of logical qubits is not limited to $K^{(L)}=4^{L}$. Using $M$ code blocks encoded with the level-$L$ many-hypercube codes as fault-tolerant quantum registers [44], we can use $MK^{(L)}$ logical qubits for fault-tolerant quantum computing.
@@
We developed a high-performance decoder dedicated to the many-hypercube codes based on level-by-level minimum distance decoding. (Our proposed decoding method will also be useful for other concatenated codes, such as the above concatenated quantum Hamming codes.) Using this decoder, we achieved a threshold of 5.6% for bit-flip errors, which is comparable to the surface code (10.9%) [47] and a 4%-rate qLDPC (hypergraph product) code (7.5%) [33]. We also propose fault-tolerant zero-state encoders for the many-hypercube codes. Using them, we achieved a threshold of 0.9% for a logical controlled-NOT (CNOT) gate in a circuit-level noise model. Finally, we explain how to perform logical gates for the many-hypercube codes in parallel.
II Many-hypercube codes
We start with the definition of the ${[\![6,4,2]\!]}$ code. The ${[\![6,4,2]\!]}$ code is one of the simplest stabilizer codes [3] and has only two stabilizer generators (check operators): ${S_{Z}=Z_{1}Z_{2}Z_{3}Z_{4}Z_{5}Z_{6}}$ and ${S_{X}=X_{1}X_{2}X_{3}X_{4}X_{5}X_{6}}$, which can detect an $X$ (bit-flip) error and a $Z$ (phase-flip) error, respectively (see Appendix A for the definitions of elementary gates). Its four logical qubits are defined by the following logical $Z$ and $X$ operators:
$Z_{L1}$ $=Z_{1}Z_{2},~{}X_{L1}=X_{2}X_{3},$ (1)
$Z_{L2}$ $=Z_{2}Z_{3},~{}X_{L2}=X_{1}X_{2},$ (2)
$Z_{L3}$ $=Z_{4}Z_{5},~{}X_{L3}=X_{5}X_{6},$ (3)
$Z_{L4}$ $=Z_{5}Z_{6},~{}X_{L4}=X_{4}X_{5}.$ (4)
Other definitions are possible [2], but we use this because of the geometrical interpretation of the code structure using hypercubes explained below. By the definition, the logical all-zero state of the ${[\![6,4,2]\!]}$ code is the six-qubit Greenberger-Horne-Zeilinger (GHZ) state:
$|0000\rangle_{L}=\frac{|000000\rangle+|111111\rangle}{\sqrt{2}}.$ (5)
Therefore, the zero-state encoder of the ${[\![6,4,2]\!]}$ code is given by, e.g., the quantum circuit shown in Fig. 1A. This can be generalized to an arbitrary-state encoder by adding two CNOT gates, as shown in Fig. 1B. The logical SWAP gates between the logical qubits 1 and 2, 3 and 4, and (1,2) and (3,4) can also be performed by physical SWAP gates between the physical qubits 1 and 3, 4 and 6, and (1,2,3) and (4,5,6), respectively, as shown in Fig. 1C. We can also easily find that transversal logical CNOT and Hadamard gates can be performed by transversal physical CNOT and Hadamard (and SWAP) gates, respectively, as shown in Figs. 1D and 1E.
Figure 1: Encoders and logical gates of the ${[\![6,4,2]\!]}$ code. (A) Zero-state encoder. (B) Arbitrary-state encoder. In (B), additional two CNOT gates are highlighted in red. (C) Logical SWAP gates. (D) Transversal logical CNOT gates. (E) Transversal logical Hadamard gates. In (C–E), the left and right sides correspond to logical and physical ones, respectively.
@@ II Many-hypercube codes
Figure 2: Visualization of the structure of the level-3 many-hypercube code. The vertices correspond to physical qubits. Examples of logical $Z$ and $X$ operators and stabilizers in Eqs. (12)–(19) are highlighted by blue ($Z$) and red ($X$). In general, logical $Z$ and $X$ operators for the level-$L$ many-hypercube code correspond to $L$-dimensional hypercubes, the number of which is equal to that of the logical qubits. Also, level-$l$ stabilizers correspond to $l$-dimensional objects on the hypercubes.
@@
Using the ${[\![6,4,2]\!]}$ code recursively, e.g., the level-3 many-hypercube code is defined as follows. Each of physical and encoded qubits is labelled by three integers (corresponding to the three levels) and denoted by Q with subscripts of the integers. The corresponding operators are also labelled similarly. The four level-1 qubits $\{\mathrm{Q}_{i^{\prime},j,k}^{(1)}|i^{\prime}=1,\ldots,4\}$ are encoded into the corresponding six physical qubits $\{\mathrm{Q}_{i,j,k}|i=1,\ldots,6\}$ with the ${[\![6,4,2]\!]}$ code ($j,k=1,\ldots,6$), where the two stabilizers and encoded $Z$ and $X$ operators are defined as follows:
$SZ_{j,k}^{(1)}$ $=\bigotimes_{i=1}^{6}Z_{i,j,k},$ (6)
$SX_{j,k}^{(1)}$ $=\bigotimes_{i=1}^{6}X_{i,j,k},$ (7)
$Z_{1,j,k}^{(1)}$ $=Z_{1,j,k}Z_{2,j,k},~{}X_{1,j,k}^{(1)}=X_{2,j,k}X_{3,j,k},$ (8)
$Z_{2,j,k}^{(1)}$ $=Z_{2,j,k}Z_{3,j,k},~{}X_{2,j,k}^{(1)}=X_{1,j,k}X_{2,j,k},$ (9)
$Z_{3,j,k}^{(1)}$ $=Z_{4,j,k}Z_{5,j,k},~{}X_{3,j,k}^{(1)}=X_{5,j,k}X_{6,j,k},$ (10)
$Z_{4,j,k}^{(1)}$ $=Z_{5,j,k}Z_{6,j,k},~{}X_{4,j,k}^{(1)}=X_{4,j,k}X_{5,j,k}.$ (11)
@@
Similarly, the four level-2 qubits $\{\mathrm{Q}_{i^{\prime},j^{\prime},k}^{(2)}|j^{\prime}=1,\ldots,4\}$ are encoded into the corresponding six level-1 qubits $\{\mathrm{Q}_{i^{\prime},j,k}^{(1)}|j=1,\ldots,6\}$ ($i^{\prime}=1,\ldots,4,k=1,\ldots,6$) and the four level-3 qubits $\{\mathrm{Q}_{i^{\prime},j^{\prime},k^{\prime}}^{(3)}|k^{\prime}=1,\ldots,4\}$ are encoded into the corresponding six level-2 qubits $\{\mathrm{Q}_{i^{\prime},j^{\prime},k}^{(2)}|k=1,\ldots,6\}$ ($i^{\prime},j^{\prime}=1,\ldots,4$). In this case, we use the 64 level-3 encoded qubits as logical qubits in a fault-tolerant quantum register for quantum computing. In general, the level-$L$ many-hypercube code is defined similarly.
Interestingly, the above code structure can be visualized as shown in Fig. 2 by placing the physical qubit $\mathrm{Q}_{i,j,k}$ at the point with the coordinates ${(x,y,z)=(i,j,k)}$ in three-dimensional space, where the logical $Z$ and $X$ operators correspond to cubes. For example,
$Z_{1,1,1}^{(3)}$ $=Z_{1,1,1}^{(2)}Z_{1,1,2}^{(2)}$
$=Z_{1,1,1}^{(1)}Z_{1,2,1}^{(1)}Z_{1,1,2}^{(1)}Z_{1,2,2}^{(1)}$
$=Z_{1,1,1}Z_{2,1,1}Z_{1,2,1}Z_{2,2,1}Z_{1,1,2}Z_{2,1,2}Z_{1,2,2}Z_{2,2,2}$ (12)
$X_{1,1,1}^{(3)}$ $=X_{1,1,2}^{(2)}X_{1,1,3}^{(2)}$
$=X_{1,2,2}^{(1)}X_{1,3,2}^{(1)}X_{1,2,3}^{(1)}X_{1,3,3}^{(1)}$
$=X_{2,2,2}X_{3,2,2}X_{2,3,2}X_{3,3,2}X_{2,2,3}X_{3,2,3}X_{2,3,3}X_{3,3,3}$ (13)
are highlighted in blue and red, respectively, in Fig. 2. Also, the level-$l$ stabilizers correspond to $l$-dimensional objects on the cubes, e.g., as follows:
$SZ_{1,4}^{(1)}$ $=\bigotimes_{i=1}^{6}Z_{i,1,4},$ (14)
$SX_{1,5}^{(1)}$ $=\bigotimes_{i=1}^{6}X_{i,1,5},$ (15)
$SZ_{1,6}^{(2)}$ $=\bigotimes_{j=1}^{6}Z_{1,j,6}^{(1)}=\bigotimes_{j=1}^{6}Z_{1,j,6}Z_{2,j,6},$ (16)
$SX_{1,6}^{(2)}$ $=\bigotimes_{j=1}^{6}X_{1,j,6}^{(1)}=\bigotimes_{j=1}^{6}X_{2,j,6}X_{3,j,6},$ (17)
$SZ_{3,1}^{(3)}$ $=\bigotimes_{k=1}^{6}Z_{3,1,k}^{(2)}=\bigotimes_{k=1}^{6}Z_{3,1,k}^{(1)}Z_{3,2,k}^{(1)}$
$=\bigotimes_{k=1}^{6}Z_{4,1,k}Z_{5,1,k}Z_{4,2,k}Z_{5,2,k},$ (18)
$SX_{3,2}^{(3)}$ $=\bigotimes_{k=1}^{6}X_{3,2,k}^{(2)}=\bigotimes_{k=1}^{6}X_{3,1,k}^{(1)}X_{3,2,k}^{(1)}$
$=\bigotimes_{k=1}^{6}X_{5,1,k}X_{6,1,k}X_{5,2,k}X_{6,2,k}.$ (19)
which are also highlighted in blue ($SZ$) and red ($SX$) in Fig. 2.
In general, the logical $Z$ and $X$ operators of the level-$L$ many-hypercube code correspond to $L$-dimensional hypercubes in $L$-dimensional space and also its level-$l$ stabilizers correspond to $l$-dimensional objects on the hypercubes, where the number of the hypercubes is equal to that of the logical qubits. This is the reason for the name of the proposed code. [The “many” is used to distinguish it from a conventional hypercube code [48], which is the ${[\![2^{D},D,2]\!]}$ code defined on a single $D$-dimensional hypercube.]
@@
The number of the vertices of each $L$-dimensional hypercube is $2^{L}$, which is equal to the code distance. That is, the size of the many-hypercube codes is increased by increasing not the edge length but the dimension while keeping the edge length. In contrast, the logical operators and stabilizers of topological codes, such as the surface code [4, 5, 6, 7, 8, 9, 10] and the color code [49, 49, 50, 52], correspond to the edges and faces, respectively, and their code sizes are increased by increasing the edge length. Unlike the topological codes, the many-hypercube codes require interactions beyond nearest-neighbor ones. This is experimentally challenging, but recent experimental advances in ion-trap [53, 53, 54, 55, 56, 57, 58, 59] and neutral-atom [60, 61, 62, 63] systems are encouraging. In fact, hypercube connectivity has already been realized experimentally [63].
III Decoders